OpenStudy (anonymous):

solve by using a table: x^2 - 6x + 8 = 0

OpenStudy (anonymous):

what table?

OpenStudy (anonymous):

it's a simple quadratic equation

OpenStudy (anonymous):

D=b^2-4*a*c D=36-32=4 x1=(6+sqrt(4))/2=4 x2=(6-sqrt(4))/2=2

OpenStudy (mathmate):

Let y=x^2-6x+8 Make the following table x y 0 8 1 3 2 0 3 -1 4 0 5 3 .... Locate the roots of equation y=x^2-6x+8=0

OpenStudy (anonymous):

Mathmate, Im not sure how to do it can you walk me through it.

OpenStudy (mathmate):

The equation you are trying to solve is y=x^2-6x+8=0 So we need to find the values of x such that y=0. The table is x against y, which means that when x=0, you would substitute 0 for x, such as y=0^2-6(0)+8=0-0+8=8. For x=1, we do the same, y=1^2-6(1)+8=1-6+8=3 We continue and complete the table until we find two zeroes (at x=2 and x=4). This means that when x=2 or when x=4, y=x^2-6x+8=0, which is precisely the solution we're looking for, the solution of the given equation.

OpenStudy (anonymous):

oh okay, thank you so much for the explanation

OpenStudy (mathmate):

My pleasure! :)